# Using Dynamic Safety Margins as Control Barrier Functions

### v4

@@ 6.2 Overhead Crane
$M(\boldsymbol{\mathbf{q}})\ddot{\boldsymbol{\mathbf{q}}}+C(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})\dot{\boldsymbol{\mathbf{q}}}+G(\boldsymbol{\mathbf{q}})=Bu,$ (49)
where the generalized coordinates $\boldsymbol{\mathbf{q}}=[x;~\theta]$ are the gantry position $x$ and the payload angle $\theta$, and
$M(\boldsymbol{\mathbf{q}})$ $=\begin{bmatrix}m_{c}+m_{p}&m_{p}L\cos\theta\\ m_{p}L\cos\theta&m_{p}L^{2}\end{bmatrix},\quad B=\begin{bmatrix}1\\ 0\end{bmatrix}$,
$C(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})$ $=\begin{bmatrix}0&-m_{p}L\dot{\theta}\sin\theta\\ 0&0\end{bmatrix},\quad G(\boldsymbol{\mathbf{q}})=\begin{bmatrix}0\\ m_{p}gL\sin\theta\end{bmatrix}$,
@@
It can be shown that, letting $\boldsymbol{\mathbf{x}}=[\boldsymbol{\mathbf{q}};~\dot{\boldsymbol{\mathbf{q}}}]$, the system is control-affine and can be written as in (1). We consider the prestabilizing PD control law given in [44]
$\pi(\boldsymbol{\mathbf{x}},v)=-k_{p}(x-v)-k_{d}\dot{x},$ (50)
@@
$V(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{q}})\dot{\boldsymbol{\mathbf{q}}}+m_{p}gL(1-\cos\theta)+\frac{1}{2}k_{p}(x-v)^{2}.$
The storage function $V$ satisfies Definition 4.23 with $\mathcal{D}_{v}=\mathbb{R}\times(-\pi,\pi)\times\mathbb{R}^{2}$ and also satisfies Assumption 4. We now define four types of constraints and, for each one, we find the associated safety threshold value or a continuously differentiable bounding function. For details on the construction of each $\Gamma^{*}_{i}(v)$, the reader is referred to [32, 42].
@@
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [32, Proposition 4], we can define
    $\underline{V}(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{x}})\dot{\boldsymbol{\mathbf{q}}}+\frac{4}{\pi^{2}}m_{p}gL\theta^{2}+\frac{1}{2}k_{p}(x-v)^{2},$
    which satisfies $\underline{V}(\boldsymbol{\mathbf{x}},v)\leq V(\boldsymbol{\mathbf{x}},v)$, $\forall\theta\in(-\pi/2,\pi/2)$. Moreover, since $x_{\text{max}}\leq p_{\text{max}}$, we can perform a planar embedding of the payload constraint
@@
The stability threshold value is
$\overline{\Gamma}(v)=2m_{p}gL.$
Using Theorem 4.27, we conclude that
$\Delta(\boldsymbol{\mathbf{x}},v)=\begin{bmatrix}\Gamma^{*}_{1}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{2}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{3}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma^{*}_{4}(v)-V(\boldsymbol{\mathbf{x}},v)\\ \Gamma_{5}(v)-V(\boldsymbol{\mathbf{x}},v)\\ 0.99\overline{\Gamma}(v)-V(\boldsymbol{\mathbf{x}},v)\end{bmatrix}$,
is a DSM for $\pi$ and, by Theorem 3.9, a (vector-valued) CBF with respect to the augmented system. Note that the stability threshold value satisfies $\Gamma^{*}_{4}(v)<\overline{\Gamma}(v)$. As such, when the angle constraints are enforced, there is no need to include $\overline{\Gamma}(v)$ in the definition of $\Delta$ above.
@@
    Nominal: The nominal controller is a PD control law $\kappa(\boldsymbol{\mathbf{x}},r)=-k_{p,\kappa}(x-r)-k_{d,\kappa}\dot{x}$, where the gains $k_{p,\kappa}>0$ and $k_{d,\kappa}>0$ can differ from the ones in (50) and $r\in\mathbb{R}$ is the target reference.
@@
    $\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{v}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\kappa(\boldsymbol{\mathbf{x}},v)\\ \min_{i}\Delta_{i}(\boldsymbol{\mathbf{x}},v)\rho_{\text{ERG}}(v,r)\end{bmatrix}$. (52)
@@
    $\psi(\boldsymbol{\mathbf{x}})=u_{\max}\tanh\left(\frac{-K_{\psi}\boldsymbol{\mathbf{x}}}{u_{\max}}\right),$ (53)
    where $K_{\psi}=[3.15~-16.35~5.48~-5.16]$ was obtained linearizing the system around the origin and solving the LQR problem. The $\tanh$ term is a smooth clipping function that ensures $|\psi(\boldsymbol{\mathbf{x}})|\leq u_{\max}$. An invariant set associated with this backup policy is $\mathcal{S}^{0}=\{\boldsymbol{\mathbf{x}}\in\mathbb{R}^{4}\mid\boldsymbol{\mathbf{x}}^{\top}P_{\psi}\boldsymbol{\mathbf{x}}\leq\Gamma_{\psi}\}$, where $P_{\psi}$ is the symmetric, positive definite matrix that solves the Riccati equation associated with the LQR problem, and $\Gamma_{\psi}=2.5$ is determined heuristically through trial and error in the simulation. We formulate the backup CBF
    $h^{\psi}(\boldsymbol{\mathbf{x}})=\min\left[\min_{\tau\in[0,T]}h^{\mathcal{C}}\big(\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\big),h^{\mathcal{S}}\big(\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},T)\big)\right],$
    where $T$ is the simulation horizon, $\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}}_{0},\tau)$ is the solution of the backup dynamics at time $\tau$ with initial condition $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$,
    $h^{\mathcal{C}}(\boldsymbol{\mathbf{x}})=\min\begin{bmatrix}x_{\max}-|x|\\ \theta_{\max}-|\theta|\\ p_{\max}-x-L\sin\theta\end{bmatrix}$, (54)
    and $h^{\mathcal{S}}(\boldsymbol{\mathbf{x}})=\Gamma_{\psi}-\boldsymbol{\mathbf{x}}^{\top}P_{\psi}\boldsymbol{\mathbf{x}}$. The class $\mathscr{K}$ functions used are $\alpha(c)=40c$ for the position and payload constraints, $\alpha(c)=100c$ for the angle constraints and $\alpha(c)=400c$ for the terminal constraint $h^{\mathcal{S}}$.
Table 2: Overhead Crane Simulation Parameters
@@
General parameters for the example are provided in Table 2, the performance of each approach is shown in Figure 5, and the computational burden of each approach is shown in Table 1. While all approaches achieve the goal and respect the safety constraints, their performance differs significantly. The ERG exhibits the slowest response but has the least computational burden. On the other hand, the backup CBF approach tracks the nominal behavior closely at the expense of high computational cost. Our DSM-CBF approach sits in the middle with a strict performance improvement over the ERG at a modest computational burden.
Figure 5: Closed-loop behavior of the overhead crane system under each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and DSM-CBF approaches, respectively. While only the trace for horizon $T=5$ for the backup CBF approach is shown, the computational burden of other horizon choices is shown in Table 1.
7 Conclusion

### v5

@@ VI-B Overhead Crane
$M(\bm{q})\ddot{\bm{q}}+C(\bm{q},\dot{\bm{q}})\dot{\bm{q}}+G(\bm{q})=Bu,$ (46)
where the generalized coordinates $\bm{q}=[x;~\theta]$ are the gantry position $x$ and the payload angle $\theta$, and
$M(\bm{q})$ $=\begin{bmatrix}m_{c}+m_{p}&m_{p}L\cos\theta\\ m_{p}L\cos\theta&m_{p}L^{2}\end{bmatrix},\quad B=\begin{bmatrix}1\\ 0\end{bmatrix}$,
$C(\bm{q},\dot{\bm{q}})$ $=\begin{bmatrix}0&-m_{p}L\dot{\theta}\sin\theta\\ 0&0\end{bmatrix},\quad G(\bm{q})=\begin{bmatrix}0\\ m_{p}gL\sin\theta\end{bmatrix}$,
@@
It can be shown that, letting $\bm{x}=[\bm{q};~\dot{\bm{q}}]$, the system is control-affine and can be written as in (1). We consider the prestabilizing PD control law given in [19]
$\pi(\bm{x},v)=-k_{p}(x-v)-k_{d}\dot{x},$ (47)
@@
$V(\bm{x},v)=\frac{1}{2}\dot{\bm{q}}^{\top}M(\bm{q})\dot{\bm{q}}+m_{p}gL(1-\cos\theta)+\frac{1}{2}k_{p}(x-v)^{2}.$
The storage function $V$ satisfies Definition 4 with $\mathcal{D}_{v}=\mathbb{R}\times(-\pi,\pi)\times\mathbb{R}^{2}$ and also satisfies Assumption 4. We now define four types of constraints and, for each one, we find the associated safety threshold value or a continuously differentiable bounding function. For details on the construction of each $\Gamma^{*}_{i}(v)$, the reader is referred to [37, 20].
@@
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [37, Proposition 4], we can define
    $\underline{V}(\bm{x},v)=\frac{1}{2}\dot{\bm{q}}^{\top}M(\bm{x})\dot{\bm{q}}+\frac{4}{\pi^{2}}m_{p}gL\theta^{2}+\frac{1}{2}k_{p}(x-v)^{2},$
    which satisfies $\underline{V}(\bm{x},v)\leq V(\bm{x},v)$, $\forall\theta\in(-\pi/2,\pi/2)$. Moreover, since $x_{\text{max}}\leq p_{\text{max}}$, we can perform a planar embedding of the payload constraint
@@
The stability threshold value is $\overline{\Gamma}(v)=2m_{p}gL$. Using Theorem 7, we conclude that
$\Delta(\bm{x},v)=\begin{bmatrix}\Gamma^{*}_{1}(v)-V(\bm{x},v)\\ \Gamma^{*}_{2}(v)-V(\bm{x},v)\\ \Gamma^{*}_{3}(v)-V(\bm{x},v)\\ \Gamma^{*}_{4}(v)-V(\bm{x},v)\\ \Gamma_{5}(v)-V(\bm{x},v)\\ 0.99\overline{\Gamma}(v)-V(\bm{x},v)\end{bmatrix}$,
is a DSM and, by Theorem 4, a CBF with respect to the augmented system. Note that the stability threshold value satisfies $\Gamma^{*}_{4}(v)<\overline{\Gamma}(v)$. As such, when the angle constraints are enforced, there is no need to include $\overline{\Gamma}(v)$ in the definition of $\Delta$ above.
@@
    Nominal: The nominal controller is a PD control law $\kappa(\bm{x},r)=-k_{p,\kappa}(x-r)-k_{d,\kappa}\dot{x}$, where the gains $k_{p,\kappa}>0$ and $k_{d,\kappa}>0$ can differ from the ones in (47) and $r\in\mathbb{R}$ is the target reference.
@@
    $\begin{bmatrix}\dot{\bm{x}}\\ \dot{v}\end{bmatrix}=\begin{bmatrix}f(\bm{x})+g(\bm{x})\kappa(\bm{x},v)\\ \min_{i}\Delta_{i}(\bm{x},v)\rho_{\text{ERG}}(v,r)\end{bmatrix}$. (49)
@@
    $\psi(\bm{x})=u_{\max}\tanh\left(\frac{-K_{\psi}\bm{x}}{u_{\max}}\right),$ (50)
    where $K_{\psi}=[3.15~-16.35~5.48~-5.16]$ was obtained linearizing the system around the origin and solving the LQR problem. The $\tanh$ term is a smooth clipping function that ensures $|\psi(\bm{x})|\leq u_{\max}$. An invariant set associated with this backup policy is $\mathcal{S}^{0}=\{\bm{x}\in\mathbb{R}^{4}\mid\bm{x}^{\top}P_{\psi}\bm{x}\leq\Gamma_{\psi}\}$, where $P_{\psi}$ is the symmetric, positive definite matrix that solves the Riccati equation associated with the LQR problem, and $\Gamma_{\psi}=2.5$ is determined heuristically through trial and error in the simulation. We formulate the backup CBF
    $h^{\psi}(\bm{x})=\min\left[\min_{\tau\in[0,T]}h^{\mathcal{C}}\big(\Phi_{f_{\psi}}(\bm{x},\tau)\big),h^{\mathcal{S}}\big(\Phi_{f_{\psi}}(\bm{x},T)\big)\right],$
    where $T$ is the simulation horizon, $\Phi_{f_{\psi}}(\bm{x}_{0},\tau)$ is the solution of the backup dynamics at time $\tau$ with initial condition $\bm{x}(0)=\bm{x}_{0}$,
    $h^{\mathcal{C}}(\bm{x})=\min\begin{bmatrix}x_{\max}-|x|\\ \theta_{\max}-|\theta|\\ p_{\max}-x-L\sin\theta\end{bmatrix}$, (51)
    and $h^{\mathcal{S}}(\bm{x})=\Gamma_{\psi}-\bm{x}^{\top}P_{\psi}\bm{x}$. The class $\mathscr{K}$ functions used are $\alpha(c)=40c$ for the position and payload constraints, $\alpha(c)=100c$ for the angle constraints and $\alpha(c)=400c$ for the terminal constraint $h^{\mathcal{S}}$.
TABLE II: Overhead Crane Simulation Parameters
@@
General parameters for the example are provided in Table II, the performance of each approach is shown in Figure 5, and the computational burden of each approach is shown in Table I. While all approaches achieve the goal and respect the safety constraints, their performance differs significantly. The ERG exhibits the slowest response but has the least computational burden. On the other hand, the backup CBF approach tracks the nominal behavior closely at the expense of high computational cost. Our DSM-CBF approach sits in the middle with a strict performance improvement over the ERG at a modest computational burden.
Figure 5: Closed-loop behavior of the overhead crane system under each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and DSM-CBF approaches, respectively. While only the trace for horizon $T=5$ for the backup CBF approach is shown, the computational burden of other horizon choices is shown in Table I.
VII Conclusion